### v1

# Quasinormal modes of spontaneous scalarized Kerr black holes

Abstract
Recent studies reveal that Kerr black holes can undergo spontaneous scalarization via an appropriate coupling between a scalar field and the Gauss-Bonnet term. This mechanism introduces observable signatures in the quasinormal mode spectrum, which characterizes the ringdown phase of gravitational waves. In this work, we present, for the first time, the quasinormal modes of rotating scalarized black holes within the framework of Einstein-scalar-Gauss-Bonnet theory, without resorting to perturbative expansion of either the coupling constant or the black hole spin. Using spectral decomposition, we numerically construct both the background solution and the perturbation functions, reformulating the calculation of quasinormal modes as a generalized eigenvalue problem. Our results pave the way for advanced ringdown tests of spontaneous scalarization using gravitational wave data.
Introduction — The emerging field of gravitational wave (GW) astronomy offers an unprecedented opportunity to explore fundamental physics in strong gravity regime LIGOScientific:2018mvr . As GW data grows and detector sensitivity improves, General Relativity (GR) – the most successful theory of gravity – can be tested from various perspectives Berti:2015itd . One of GR’s most remarkable predictions, black holes (BHs), plays a key role in probing strong gravity due to their simplicity Barack:2018yly . According to the no-hair theorem Herdeiro:2015waa , BHs in GR are uniquely characterized by mass, spin, and electric charge. Current and future GW observations enable precise tests of this theorem, including the potential existence of extra scalar fields or other matter fields that hint at new physics Isi:2019aib ; Barausse:2020rsu ; Brito:2017wnc ; Hannuksela:2018izj .
Spontaneous scalarization is an intriguing mechanism that introduces an extra scalar degree of freedom for BHs through the coupling of a scalar field with a source term, such as curvature or matter Antoniou:2017acq ; Doneva:2017bvd ; Antoniou:2017hxj ; Doneva:2018rou ; Collodel:2019kkx ; Herdeiro:2020wei ; Silva:2017uqg ; Berti:2020kgk ; Herdeiro:2018wub ; Silva:2017uqg ; Cunha:2019dwb . The spontaneous scalarization model can withstand rigorous tests in weak gravity regime wherein this model is indistinguishable from GR. On the other hand, this model allows a sharp transition to a stationary scalarized BH that can differ significantly from GR in the strong field regime, and hence the GWs emitted by the scalarized BHs may possess detectable deviations from GR expectations. For recent progress on spontaneous scalarization, one can refer to the review Doneva:2022ewd .
Recently, the Einstein-scalar-Gauss-Bonnet (EsGB) theory becomes the most attractive model in which spontaneous scalarization can occur Silva:2017uqg ; Berti:2020kgk ; Herdeiro:2020wei ; Antoniou:2017acq ; Doneva:2017bvd ; Antoniou:2017hxj ; Doneva:2018rou ; Collodel:2019kkx ; Herdeiro:2020wei ; Cunha:2019dwb . In this theory, the scalar field couples to the GB invariant, a higher-order curvature term motivated by high-energy physics Zwiebach:1985uq , which also avoids Ostrogradsky instabilities Ostrogradsky:1850fid . Spontaneous scalarization occurs when the vacuum solution, represented by the Kerr BH, undergoes a tachyonic instability under scalar perturbations as the model parameters exceed a certain threshold Doneva:2020nbb ; Dima:2020yac , leading to the dynamic formation of a stationary scalar field profile near the event horizon East:2021bqk . As a viable model that mimic the Kerr phenomenology, the scalarized BHs in EsGB theory have been studied in different aspects, including the BH shadow Cunha:2019dwb , the dynamical evolution East:2021bqk and the stability analysis Blazquez-Salcedo:2020rhf ; Blazquez-Salcedo:2020caw ; Blazquez-Salcedo:2018jnn ; Silva:2018qhn ; Macedo:2020tbm .
A straightforward approach to test the EsGB theory is measuring the effects of scalarization on GWs of coalescing NSs or BHs. The constraint on EsGB theory with inspiral GW data by considering the absence of scalar dipole radiation has been performed in Wong:2022wni . Nonetheless, the waveform models for the merger and ringdown stages remain underdeveloped. The ringdown waveform can be precisely modeled by the superposition of damped sinusoids with complex eigenfrequencies known as quasinormal modes (QNMs) Kokkotas:1999bd ; Berti:2009kk ; Konoplya:2011qq . The real and imaginary part of each mode, which represents the oscillation frequency and the inverse of the damping time, uniquely determined by the mass and angular momentum of the BH remnant Berti:2005ys .
To perform ringdown tests, the QNM spectrum of rotating scalarized BHs must be determined in advance. The QNMs in EsGB theory have been calculated, but only within the static limit Blazquez-Salcedo:2020rhf ; Blazquez-Salcedo:2020caw . However, calculating QNMs for a general stationary axisymmetric BH remains a recognized challenge. The BH background is no longer classified as Petrov type D; hence, the underlying assumptions of the Teukolsky formalism Teukolsky:1973ha are invalid, and it cannot be directly applied without modifications. In recent years, significant efforts have been devoted to deriving perturbation equations and the calculation of QNMs for rotating BHs within the framework of modified gravity theories, including modified Teukolsky formalism Li:2022pcy ; Hussain:2022ins , full numerical relativity simulations Okounkova:2019zjf ; AresteSalo:2022hua and small-spin approximations Wagle:2021tam ; Pierini:2022eim . In particular, a novel method relying on spectral decomposition has been introduced, through which highly accurate QNMs can be directly obtained by solving the equations governing metric and scalar field perturbations Chung:2023zdq ; Chung:2023wkd ; Blazquez-Salcedo:2023hwg ; Chung:2024ira ; Blazquez-Salcedo:2024oek .
In this paper, we study the QNM spectrum of rotating scalarized BHs in the EsGB theory. Our calculations are based on fully non-perturbative methods, where both the background solutions and perturbation functions are numerically constructed by the spectral decomposition. The corresponding dissipative boundary conditions at the event horizon or the radial infinity are presented in a compact form for the first time. We concentrate on the $l=m=2$ fundamental mode, which is the primary excitation of the ringdown phase for nearly equal-mass binary mergers validated by numerical relativity Buonanno:2006ui ; Berti:2007fi .
The EsGB model — The EsGB theory is described by the action
$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}[R-2\nabla_{\mu}\nabla^{\mu}\phi+\lambda^{2}f(\phi)\mathcal{R}_{\textrm{GB}}],$ (1)
where the GB term $\mathcal{R}_{\textrm{GB}}\equiv R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ is coupled to a function $f(\phi)$ of the scalar field $\phi$ via the coupling constant $\lambda$. The equations of motion for the backgroud and the perturbation field can be expressed as
00 0 $=$ $\mathcal{E}_{\mu\nu}^{(0)}+\epsilon\mathcal{E}_{\mu\nu}^{(1)}+\mathcal{O}(\epsilon)^{2},$
00 0 $=$ $\mathcal{S}^{(0)}+\epsilon\mathcal{S}^{(1)}+\mathcal{O}(\epsilon)^{2},$ (2)
where $\mathcal{E}_{\mu\nu}$ is the tensor of modified Einstein equation and $\mathcal{S}$ represents the scalar field equation. The superscripts $(0)$ and $(1)$ mark the equations for the background fields and the first-order perturbed fields, respectively. The coupling function $f(\phi)$ for the spontaneous scalarization model typically satisfies the conditions $f^{\prime}(0)=0$ and $f^{\prime\prime}(0)\neq 0$. We adopt the general form of $f(\phi)$ as
$f(\phi)=\frac{1}{12}(1-e^{-6\phi^{2}}),$ (3)
by following Doneva:2017bvd ; Cunha:2019dwb ; Wong:2022wni . The existence domain Cunha:2019dwb ; Wong:2022wni indicates that BHs with moderate spin ($\chi\approx 0.6$, $\chi$ is the dimensionless spin of BHs) or lower must undergo scalarization if their characteristic scale is comparable to $\lambda$. In contrast, rapidly rotating BHs remain effectively indistinguishable from Kerr BHs predicted by GR. The coupling function with the form of (3) is known as the curvature-induced scalarization. Another model called spin-induced scalarization predicts scalarization exclusively for rapidly rotating BHs, achieved by simply multiplying the coupling function in (3) by $-1$ Herdeiro:2020wei ; Berti:2020kgk .
The numerical method
Figure 1: The migration of the $l=m=2$ fundamental mode for scalarized static BHs on the complex plane with increasing $\lambda/M$. The crosses denote the modes for the gravitational perturbations (black) and the scalar perturbation (green) of the vacuum Schwarzschild BH at $\lambda=0$, respectively. As $\lambda/M$ increases, the tensor mode remains unchanged and the scalar mode varies, as depicted by the green solid line. Scalarized BHs emerge at the threshold $\lambda_{0}/M\approx 1.704$.
The metric ansatz for the background spontaneous scalarized Kerr BHs can be written as Cunha:2019dwb ; Herdeiro:2020wei
$ds^{2}$ $=$ $-e^{2F_{0}}Ndt^{2}+e^{2F_{1}}\left(dr^{2}/N+r^{2}d\theta^{2}\right)$ (4)
$+e^{2F_{2}}r^{2}\sin^{2}\theta(d\varphi-W/r^{2}dt)^{2},$
with $N\equiv 1-r_{H}/r$ and the event horizon radius $r_{H}$. The components ($F_{0},F_{1},F_{2},W$) together with the backgroud scalar field $\phi^{(0)}$, depend on $r$ and $\theta$ only, are decomposed through a spectral method developed by Fernandes:2022gde . The large system of algebraic equations for the coefficients, generated from substituting the spectral decomposition into the backgroud equations $\mathcal{E}_{\mu\nu}^{(0)}=\mathcal{S}^{(0)}=0$, is solved by the Newton-Raphson method with an appropriate seed. We recommend interested readers to refer to Fernandes:2022gde for further details. The generic perturbations of the scalarized BHs can be separated into axial-led(A) and polar-led (B) part Blazquez-Salcedo:2023hwg
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+e^{-i\omega t+im\varphi}\epsilon(h_{\mu\nu}^{\textrm{(A)}}+h_{\mu\nu}^{\textrm{(P)}}),$ (5)
where the separation of variables $e^{-i\omega t+im\varphi}$ with frequency $\omega=\omega_{R}+\omega_{I}i$ and azimuth number $m$ is generated from the symmetries of the backgroud spacetime. The axial and polar perturbations are, respectively, given by
$h_{\mu\nu}^{\textrm{(A)}}=\left(\begin{array}[]{cccc}0&0&-imh_{0}/\sin\theta&\sin\theta\partial_{\theta}h_{0}\\ 0&0&-imh_{1}/\sin\theta&\sin\theta\partial_{\theta}h_{1}\\ -imh_{0}/\sin\theta&\sin\theta\partial_{\theta}h_{0}&0&0\\ -imh_{1}/\sin\theta&\sin\theta\partial_{\theta}h_{1}&0&0\end{array}\right),$ (6)
and
$h_{\mu\nu}^{\textrm{(P)}}=\left(\begin{array}[]{cccc}H_{0}e^{-2F_{1}}N&H_{1}&0&0\\ H_{1}&H_{2}e^{2F_{1}}/N&0&0\\ 0&0&Ke^{2F_{1}}r^{2}&0\\ 0&0&0&Ke^{2F_{2}}r^{2}\sin^{2}\theta\end{array}\right).$ (7)
The components ($h_{0},h_{1},H_{0},H_{1},H_{2},K$) for the perturbed metric are straightforwardly extended from the Regge-Wheeler gauge and the Zerilli gauge. However, the axial and polar functions do not decouple from each other in a general axisymmetric spaceetime.
Refer to caption Refer to caption
Figure 2: The spectrum of the fundamental $l=|m|=2$ QNMs for scalarized Kerr BHs are shown while varying BH spin $\chi$ and fixing the coupling constant $\lambda/M=1.75$. The two gray lines depict the prograde mode ($m=2$) with larger imaginary part and the retrograde mode ($m=-2$) with lesser imaginary part for Kerr BHs.
We substitute the perturbed metric (6) and (7) and the perturbed scalar field $\phi^{(1)}$ into the first-order perturbation equations $\mathcal{E}_{\mu\nu}^{(1)}=\mathcal{S}^{(1)}=0$, resulting in a set of seven coupled linear partial differential equations for seven perturbation components. The corresponding boundary conditions for the perturbation components can be written in a compact form
$f(r,\theta)\sim\left\{\begin{array}[]{cll}&(r-r_{H})^{-\frac{i(\omega-m\Omega_{H})}{2\kappa_{H}}},&r\rightarrow r_{H},\\ &&\\ &e^{i\omega r}r^{2iM\omega},&r\rightarrow+\infty,\end{array}\right.$ (8)
where $f(r,\theta)$ represents an arbitrary component of ($h_{0},h_{1},H_{0},H_{1},H_{2},K,\phi^{(1)}$), $\Omega_{H}\equiv\frac{W}{r^{2}}|_{r=r_{H}}$ is the angular velocity at the event horizon and $\kappa_{H}\equiv\frac{1}{2r_{H}}e^{F_{0}-F_{1}}|_{r=r_{H}}$ is the surface gravity. The boundary condition at the event horizon can be derived by ensuring the regularity of perturbations there Teukolsky:1973ha ; Xiong:2024 , and can be verified through the perturbation equation of a test scalar field as suggested by Pierini:2021jxd . These boundary conditions are consistent with those of the vacuum Kerr BHs and the static scalarized BHs Blazquez-Salcedo:2020rhf .
To discretize the perturbation equations, we decompose the perturbation functions as
$f(x,\theta)=\sum^{N_{x}-1}_{i=0}\sum^{N_{\theta}-1}_{j=0}c_{ij}\ C^{(x)}_{i}(x)\ C^{(\theta)}_{j}(\theta).$ (9)
The cardinal function for $x$ is constructed by the linear combinations of Chebyshev polynomials
$C^{(x)}_{i}(x)=\sum_{n=0}^{N_{x}-1}\frac{T_{n}(x_{i})T_{n}(x)}{P^{(x)}_{n}},$ (10)
where $P^{(x)}_{n}=N_{x}$ for $n=0$ and $P^{(x)}_{n}=N_{x}/2$ for $n\neq 0$. $x\equiv 1-2r_{H}/r$ is defined by the compactification of the radial coordinate. The cardinal function for $\theta$ is
$C^{(\theta)}_{j}(\theta)=\sum_{n=0}^{N_{\theta}-1}\frac{\cos(n\ \theta_{j})\cos(n\ \theta)}{P^{(\theta)}_{n}},$ (11)
for even azimuth number $m$, with $P^{(\theta)}_{n}=N_{\theta}$ for $n=0$ and $P^{(\theta)}_{n}=N_{\theta}/2$ for $n\neq 0$. The collocation points appearing in the above cardinal functions are defined as
$x_{i}=\cos\left(\frac{(2i+1)\pi}{2N_{x}}\right),\ \ \theta_{j}=\frac{(2j+1)\pi}{2N_{\theta}},$ (12)
with $i=0,\ldots,N_{x}-1$ and $j=0,\ldots,N_{\theta}-1$. The cardinal functions (11) for the angular part differ from the Legendre polynomials, though the latter is typically the natural basis for angular components in separable perturbations of spherically symmetric BHs Blazquez-Salcedo:2023hwg . However, the two choices are equivalent, as their respective basis functions can be exactly transformed into one another.
The expansion coefficients of the spectral decomposition (9) are obtained by $c_{ij}=f(x_{i},\theta_{j})$. We utilize the co-lexicographic ordering to flatten the two-dimensional array $f(x_{i},\theta_{j})$ into a vector $\vec{f}=\{f_{I},I=1,2,...,N_{x}\times N_{\theta}\}$, and hence the differential operator acting on the perturbation function can be expressed as a Kronecker product Dias:2015nua
$\frac{d^{\alpha}}{dx^{\alpha}}\frac{d^{\beta}}{d\theta^{\beta}}f(x,\theta)\rightarrow\sum_{I=1}^{N_{x}\times N_{\theta}}f_{I}\left[[D^{(\alpha)}]^{(x)}\otimes[D^{(\beta)}]^{(\theta)}\right]_{IJ},$ (13)
with the $\alpha$-derivative matrix $[D^{(\alpha)}]^{(x)}_{ij}\equiv d^{\alpha}C^{(x)}_{i}(x_{j})/dx^{\alpha}$ and $[D^{(\alpha)}]^{(\theta)}_{ij}\equiv d^{\alpha}C^{(\theta)}_{i}(\theta_{j})/d\theta^{\alpha}$ with respect to $x$ and $\theta$, respectively.
All the steps discretize the perturbation equations $\mathcal{E}_{\mu\nu}^{(1)}=\mathcal{S}^{(1)}=0$ into a matrix equation $(\mathcal{M}_{0}+\omega\mathcal{M}_{1}+\omega^{2}\mathcal{M}_{2})\vec{f}=0$, where $\mathcal{M}_{0},\mathcal{M}_{1},\mathcal{M}_{2}$ represent rectangular matrices of order ($7\times N_{x}\times N_{\theta}$). The QNM frequency $\omega$ can be solved as a generalized eigenvalue problem for this matrix equation.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: The $l=m=2$ modes with different dimensionless coupling constant $\lambda/M$ for rotating scalarized BHs. The values of the fixed BH spin are shown at the top of each column. The axial-led, polar-led and scalar-led modes are depicted by the blue, yellow and green lines, respectively.
Results — The dynamical evolution of binary BH mergers within the framework of GR shows that the most easily excited mode is the $l=|m|=2$ fundamental mode, that is the mode with largest imaginary part Buonanno:2006ui ; Berti:2007fi . In the following, we concentrate on the $l=|m|=2$ fundamental mode.
In Fig.1, we show how the $l=|m|=2$ fundamental mode of the scalarized static BH branches from the one of the vacuum BHs and splits into three distinguishable families: axial, polar, and scalar modes, as the scaled coupling constant $\lambda/M$ (where $M$ is the mass) increases. Particularly, the scalar mode of scalarized BHs should branch from that of the vacuum BH at $\omega\approx 0.4661-0.0987i$ with the threshold $\lambda/M\approx 1.704$ instead of $\lambda=0$, as illustrated by the intersection of the green dashed and green solid lines in Fig.1. The scalar modes of vacuum BHs are calculated using the accurate continued fraction method and validated by the spectral method. The described behavior indicates that the QNMs of scalarized BHs can asymptotically approach those of vacuum BHs while $\lambda/M$ approach the threshold, analogous to the thermodynamic behavior Doneva:2017bvd . We reproduce the axial modes of the scalarized static BH reported in Blazquez-Salcedo:2020rhf . Nevertheless, in Fig.1, the polar and scalar modes exhibit discrepancies compared to Blazquez-Salcedo:2020caw . We comment that the frequency value $\omega\approx 0.481-0.0894i$ of the scalar mode of the scalarized BHs at the threshold point reported in Blazquez-Salcedo:2020caw dose not match the value of the vacuum BHs.
For scalarized Kerr BHs, the QNM modes inherit their names from those of the static limit and are named as axial, polar and scalar-led modes Blazquez-Salcedo:2024oek , even though the perturbation components are coupled in general.
Fig.2 illustrates the variation of the fundamental $l=|m|=2$ modes as the spin increases for scalarized BHs with $\lambda=1.75$, including both prograde modes ($m\omega_{R}>0$) and retrograde modes ($m\omega_{R}<0$). It is clear that the isospectrality which holds for the Kerr axial-led and polar-led modes is broken once the BHs are scalarized. We find an interesting behavior for the polar-led mode: within certain spin, the prograde mode can exhibit a smaller imaginary part compared to the retrograde mode. In contrast, the behavior of the two branches for the axial-led modes resembles that of Kerr BHs, where the prograde modes always dominate. Moreover, the real part of prograde mode shows a larger deviation from that of the Kerr BH, compared to the retrograde mode.
We demonstrate the QNMs of rotating scalarized BHs for varying coupling constant $\lambda/M$ and fixing the dimensionless spin $\chi$ in Fig.3. At low spin $\chi=0.2$, the behavior of each family is similar to that in the static BH limit as shown by Fig.1. The polar-led mode dominates only at intermediate $\lambda/M$, while the axial-led mode dominates otherwise. For larger spin $\chi=0.4$, the polar-led mode becomes dominant over larger values of the coupling constant, and the polar-led mode can even maintain dominance for $\chi=0.6$. The scalar-led mode becomes dominant only at either small or large values of the coupling constant, with the range of dominance being relatively narrow.
Discussion — QNMs of rotating BHs are crucial to the construction of ringdown waveform of binary BH mergers. In this paper, we have presented the first full calculation of the QNM spectrum of scalarized Kerr BHs in the EsGB theory, where both the BH solution and perturbations were obtained exactly. We found that the QNM with specific $l,m$ splits into three families: the axial-led, the polar-led and the scalar-led modes. The QNM for each family branch from that of vacuum BHs at the bifurcation point of $\lambda$, where the scalarized BHs are close to the vacuum BHs. We have demonstrated the behavior of $l=m=2$ fundamental modes as the BH spin or the coupling parameter $\lambda$ is varying.
The splitting of QNMs, especially the isospectrality breaking of tensor modes, presents a promising avenue for studying the ringdown phase of BHs. In GR, the $l=m=2$ fundamental mode dominates the ringdown following a merger. Assuming this dominance persists in the EsGB theory, a Bayesian analysis of the $l=m=2$ fundamental mode could uncover new physics near the BH horizon. Although detecting higher-order modes in GW signals remains challenging, this mode alone provides four independent constraints, making it a powerful probe.
Acknowledgements.
The work is in part supported by NSFC Grant No.12205104, “the Fundamental Research Funds for the Central Universities” with Grant No. 2023ZYGXZR079, the Guangzhou Science and Technology Project with Grant No. 2023A04J0651 and the startup funding of South China University of Technology.